# PHYSICISTS JOKE

### v1

@@ 5.2 Common Sense and the Universe
Original Source: New Scientist, Vol. 13, No. 271 (1962).
@@
Speaking in December 1941 at the annual meeting of the American Association for the Advancement of Science, Professor Edwin Hubble of the Mount Wilson Observatory announced that the Universe is not expanding. This was truly good news, if not for the general public, who had no reason to suspect that the Universe was expanding in the first place, then at least for those of us who humbly try to "keep up with the progress of science."
For the last twenty-five years, specifically since the day Professor de Sitter published his terrible hypothesis in 1917, we, as best we could, have tried to live in this expanding Universe, where every part flies away from every other part at a furious speed. It reminded us of that desperate lover who jumped on his horse and rode madly in all directions. The idea was majestic, but it created a certain feeling of discomfort.
@@
Nevertheless, we had to believe it. We had to rely, for example, on the authority of Spencer Jones of the Royal Astronomical Society, who in his 1940 book Life on Other Worlds claimed that "a distant galaxy in the constellation Boötes is receding from us at a speed of 24,300 miles per second." From this, it follows that it is at a distance of 230,000,000 light-years. I will remind my friends—lovers of science—that a light-year is the distance light travels in a year, moving at a speed of 186,000 miles per second. In other words, this "distant galaxy" is located away from us at a distance of 1,049,970,980,000,000,000,000 miles…
And now it turns out that it is not moving away at all! The astronomers did not simply assume that the Universe is expanding, but proved it by studying the behavior of the red part of the spectrum, which turned even redder from such a discovery, like the shy water in Cana of Galilee that "saw the Lord God and blushed." One of the most outstanding of our astronomers, Sir Arthur Eddington, wrote a book The Expanding Universe to bring this fact to general attention. Astronomers took the news of this universal explosion with the same calmness with which they once accepted the news of the impending heat death of the Universe; according to the second law of thermodynamics, it must eventually die of cold.
But the joy delivered to us by Professor Hubble is tempered by some doubts and reflections. Do not think that I express disbelief in science or disrespect for it (in our days this would be as monstrous as not believing in the Holy Trinity in the time of Isaac Newton). But still… Today we expand, tomorrow—we shrink: first we suffer in a curved and closed space, then this loop is loosened and dissolved completely; only just were we sentenced to agonizing death at a temperature of minus 273 degrees Celsius, in a cold that grips everything and everyone,—and suddenly it got warmer again.
So do we have the right to ask—what is the matter? Where are we? Einstein answers this question: "Nowhere, because there are no places where we could be at all." So pick up your books, follow the development of science, and wait for the next astronomical congress.
Take the history with the famous Second Law of Thermodynamics, this curse of inexorable fate, which dooms the entire Universe (or at least all life in the Universe) to death from cold. Now I recall with regret the tears I shed, heartily sympathizing with that last bunch of doomed ones who are destined to die at a temperature of 273 degrees below zero Celsius, at absolute zero, when all heat is exhausted and all molecules stop. Furnaces will not burn, matches will not light, and there will be no one to light them…
I remember when I first read about this cruel law in a "popular science" book titled Our Time is Running Out as a small boy. It was written by Richard Proctor and produced a terrifying impression. The Sun, it turns out, is cooling down and will go out completely. This was confirmed by Lord Kelvin. As a true Scotsman, he feared nothing and left the Sun and the entire Solar System only ninety million years of life.
This famous prediction was first made in 1824 by the great French physicist Sadi Carnot. He showed that all bodies in the Universe change their temperature—hot ones cool down, and cold ones heat up. Thus, they equalize their temperature. It is the same as dividing a rich inheritance equally among all poor relatives; the result will be general poverty. So the cold of world space must engulf us all eventually.
True, a ray of hope appeared when Ernest Rutherford and others discovered radioactivity. Radioactive atoms, decaying, seemed to be able to maintain fire on the Sun for quite a long time. This pleasant news meant that the Sun, on the one hand, is much younger, and on the other—much older than Lord Kelvin assumed. But all the same, this is only a delay. Eventually, scientists calculate, we will still freeze.
When in place of medieval superstitions came enlightenment, the first sciences that stood out and defined themselves were mathematics, astronomy, and physics. By the beginning of the nineteenth century, everything was put in its place; the Solar System rotated so sleepily and smoothly that Laplace managed to convince Napoleon that God, who would look after it, was generally unnecessary. Gravity worked like a clock. Chemistry, which in the time of Benjamin Franklin was merely a collection of incoherent experimental data, turned into a science after Dalton appeared and showed that any substance can be broken down into very, very small atoms, atoms combine into molecules, and everything goes according to plan. Since Faraday and Maxwell, electricity took its place in the new scientific order (it turned out that it is the same thing as magnetism).
Around 1880, it became clear that the world is perfectly explained by science. Metaphysics was still muttering something in its sleep. Theology still delivered sermons. It tried to dispute many discoveries of science, especially in geology and in the new evolutionary theory of life. But science already paid little attention to this.
Because everything was very simple. There is time and space—things too obvious to explain. There is matter, made of small solid atoms, similar to tiny grains. All this moves, obeying the law of universal gravitation. Nebulae condense into stars, stars eject planets, planets cool down, life is born on them, it develops and becomes rational, first anthropoid apes appear, then Bishop Wilberforce and, finally, Professor Huxley.
There remained a few small ambiguities, for example, the question of what Space, Matter, Time, Life, and Reason actually are. But Herbert Spencer very opportunely guessed to call them unknowable, locked them in a desk drawer, and left them there.
Everything was explained by Mechanical Determinism. There remained only this repulsive skeleton in the box of the writing desk. Yes, there was also something strange and mysterious in electricity, which was not exactly a thing, but not exactly a fiction. There was still a strange riddle about "action at a distance," and electricity only exacerbated it. How does gravity get from the Earth to the Sun? If there is nothing in space, then how does light fly to us from the Sun in eight minutes and even from Sirius—in eight years? Even the invention of "ether," a kind of universal jelly, did not relieve science of a certain unconvincingness.
And so, right on the threshold of the 20th century, the whole building began to crumble. First came the warning that not everything is smooth—the discovery of X-rays. Roentgen discovered them, and since then most physicists call them Roentgen rays. Sir William Crookes, experimenting with tubes filled with rarefied gas, opened "radiant matter" just as accidentally as Columbus opened America. The British government immediately granted Crookes a knighthood, but it was already too late. The deed was done.
Then followed the works of a whole school of radioactivity researchers. The crown of them was the work of Rutherford, who revolutionized the theory of the structure of matter. I knew Rutherford well—we were colleagues at McGill University for seven years—and I can confirm that he acted without a premeditated intention to shake the foundations of the Universe. But he did exactly that, for which he was also made a lord.
One should not confuse the works of Rutherford on nuclear physics with the theory of space and time created by Einstein. Rutherford never referred to Einstein in his life. Even when he worked at the Cavendish Laboratory and, showing black ingratitude, smashed the very atoms that made him famous, even then he needed nothing from Einstein. I once asked Rutherford (this was in 1923, Einstein’s worldwide fame was at its zenith) what he thought of the theory of relativity. "Ah, nonsense!" he answered. "We don’t need this for our work!" His biographer and admirer, Professor Eve, says that when the German physicist Wien told Rutherford that not a single Anglo-Saxon understands the theory of relativity, Rutherford replied: "Naturally, we have too much common sense."
But still, the main troubles started precisely with Einstein. In 1905, he announced that there is no absolute rest. And since then it has been gone. But only after the First World War did the reading public throw itself at Einstein, and shelves in shops began to break from books "about relativity."
@@
Einstein knocked out space and time just as Rutherford knocked out matter. The general view of the theory of relativity on space is very simple. Einstein explained to everyone that there is no such place as "here." "But I am here," you say. "Here—is just the place where I am sitting right now." But you are moving! Earth turns, and you turn with it. Together with the Earth, you move around the Sun, and together with the Sun—following the "distant galaxy," which itself rushes at a speed of 26,000 miles per second. So what is this place—"here"? How do you mark it? It all reminds one of the story about two idiots fishing. One says to the other: "Listen, we should mark the place where we caught this healthy fish," and the other replies: "I already made a mark on the boat." Here is your "here"!
Einstein’s discovery of the curvature of physical space was greeted with an explosion of applause, such as had not been heard perhaps only since baseball. The brilliant scientist, Sir Arthur Eddington, who handles space and time as a poet (even his discussions about gravity are permeated with humor: he says that the ideal opportunity to study gravity is available to a man falling in an elevator from the twentieth floor), so, Sir Arthur Eddington applauded louder than anyone. According to him, without this curvature, it is impossible to understand anything at all. We crawl like a fly on a globe, thinking it is flat. The mysteries of gravity puzzle us (I do not mean those few lucky ones who had the chance to fall in an elevator from the twentieth floor. But revelation descended on them too late, and the revelation consists in the following: we do not fall at all, we simply curve). "Admit the curvature of space," wrote Eddington in 1927, "and the mysterious force will disappear. Einstein banished this demon."
But now, fourteen years later, it begins to seem that Einstein is little worried whether space is curved or not. To him, apparently, it’s all the same. One famous physicist, heading a faculty at one of the largest universities, recently wrote to me on this subject: "Einstein hopes that the general theory, taking into account some properties of space resembling the fact that space is usually called curved, may in the future turn out to be more fruitful than it, apparently, takes place at the present time." Spoken purely professionally. The majority says simply that Einstein waved his hand at curved space. It is the same as if Sir Isaac Newton, yawning, said: "Ah, you are talking about this apple—maybe it didn’t fall at all?"
— It is necessary to admit, gentlemen, that this discovery is somewhat staggering.

### v2

@@ 5.2 Common Sense and the Universe
Original Source: "Last Leaves" (1945); reprinted in New Scientist (1962).
@@
Speaking in December 1941 at the annual meeting of the American Association for the Advancement of Science, Professor Edwin Hubble of the Mount Wilson Observatory made the announcement that the Universe is not expanding. This is good news.
For twenty-five years, indeed ever since Professor de Sitter launched his theory in 1917, we have been living in an expanding universe, one in which everything flies away from everything else at a terrific speed. It made us feel like the man in the story who flung himself on his horse and rode madly off in all directions. The idea was majestic, but it created a sense of dispersion.
@@
Yet we had to believe it. We had to rely on the authority of men like Spencer Jones of the Royal Astronomical Society, who told us in his Life on Other Worlds (1940) that "a distant nebula in Boötes is receding at 24,300 miles a second." We calculated that it must be 230,000,000 light years away. We knew that a light year means the distance light travels in a year at 186,000 miles a second. In other words, this "distant nebula" is 1,049,970,980,000,000,000,000 miles away.
And now it turns out it isn’t moving! The astronomers did not assume the expansion; they proved it by the behavior of the red band in the spectrum, which blushed at the idea. One of our most distinguished astronomers, Sir Arthur Eddington, wrote a book about it, The Expanding Universe. Astronomers took the explosion of the universe with the same cool detachment as when they announced that the sun is dying and that the universe will eventually freeze.
But the joy delivered to us by Professor Hubble is tempered by doubts. Do not think that I mean any disrespect to science. In our time that would be as bad as blasphemy in the time of Isaac Newton. But… Today we expand, tomorrow we contract. First we are curved, then we are straight. Just as we were going to freeze to death at absolute zero, it gets warm again.
We have a right to ask, where are we? Einstein answers: "Nowhere," because there is no place to be.
Take the Second Law of Thermodynamics, that grim decree of Fate which doomed the universe to die of cold. I look back with regret to the tears I wasted over that. I recall reading as a boy a book by Richard Proctor called Our Place Among Infinities. It was terrifying. The sun, it appeared, was cooling; Lord Kelvin proved it. He gave it a few million years. Sadi Carnot showed in 1824 that all bodies equalized their temperature—hot things cooled, cold things warmed. It was a case of dividing a rich inheritance among a lot of poor relations; the result is poverty for all.
Then came Rutherford and radioactivity. The sun, it seemed, was not cooling; it was "decaying." It was much younger than Kelvin thought, and much older. But the end was the same. We still froze.
When the medieval superstition was replaced by enlightenment, science proceeded to explain the universe. Mathematics, astronomy, and physics put everything in its place. The solar system spun so beautifully that Laplace convinced Napoleon that he didn’t need God to watch over it. Gravitation worked like a clock. Chemistry turned matter into nice little atoms, like seeds.
By 1880 the world was all explained. Metaphysics mumbled in its sleep; Theology preached sermons. But Science didn’t care. It had Space, and Time, and Matter—all genuine.
And then, at the turn of the century, the whole structure began to crumble. First came the X-rays. Roentgen discovered them. Then Sir William Crookes discovered "radiant matter" by accident. Then Rutherford knocked the bottom out of the theory of matter. I knew Rutherford well at McGill; I can testify he had no intention of destroying the universe. But he did. He showed that atoms were not hard little seeds, but holes—mostly empty space.
@@
Then came Einstein. In 1905 he announced that there is no such thing as Absolute Rest. After the Great War, the reading public seized on Einstein. He knocked out Space and Time just as Rutherford knocked out Matter. He showed that there is no "here." "But I am here," you say. "Here is where I am." But you are moving! The earth turns, the sun moves, the galaxy moves. So where is "here"? How do you mark it? It reminds one of the story of the two idiots out fishing. One said, "We should mark this place where we caught the big fish." The other said, "I’ve already marked it on the boat." There is your "here"!
Einstein’s discovery of curved space was greeted with applause. Sir Arthur Eddington, who handles space and time like a poet, applauded loudest. "Admit the curvature of space," he wrote in 1927, "and the mysterious force of gravity disappears. Einstein has exorcised the demon."
But now, fourteen years later, it seems Einstein doesn’t care if space is curved or not. A prominent physicist wrote to me: "Einstein hopes that the general theory… may in the future prove more fruitful than it seems at present." This is purely professional talk. The layman reads it as: Einstein has shrugged off curved space. It is as if Sir Isaac Newton, with a yawn, had said, "Oh, about that apple—perhaps it didn’t fall after all."
Personally, I admit, the situation is a little staggering.